# Quantum teleportation implies symmetry-protected topological order

### v2

@@ Appendix C Proof of Proposition 5
At Heisenberg time $\tau=0$, the initial operator acts as $\Gamma\otimes\widetilde{\mathds{1}}_{\rm ss}$ (i.e., trivially on all Stinespring registers). The outcome-averaged Heisenberg update to $\Gamma$ is given by
$\Gamma^{\prime}$ $=\operatorname*{tr}\limits_{{\rm ss},j}\,\left[\,\mathcal{M}^{\dagger}_{\left[A_{j}\right]}\,\Gamma\otimes\widetilde{\mathds{1}}_{j}\,\mathcal{M}_{\left[A_{j}\right]}\,\widetilde{P}^{(0)}_{j}\,\right]\,,~{}~{}$ (C.107)
@@
$\Gamma^{\prime}$ $=\sum\limits_{m,n=0,1}\frac{1}{4}\left(\mathds{1}+(-1)^{m}\,\bar{A}_{j}\right)\,\Gamma\,\left(\mathds{1}+(-1)^{m}\,\bar{A}_{j}\right)$
×$~{}~{}~{}\quad~{}~{}\times~{}{\rm tr}\left[\,\widetilde{X}^{m}_{j}\,\widetilde{\mathds{1}}_{j}\,\widetilde{X}^{n}_{j}\,\widetilde{P}^{(n)}_{j}\,\right]$
$=\frac{1}{4}\sum\limits_{m=0,1}\,\left(\Gamma+(-1)^{m}\,\left\{\bar{A}_{j},\,\Gamma\right\}+\bar{A}_{j}\,\Gamma\,\bar{A}_{j}\right)$
$=\frac{1}{4}\sum\limits_{m=0,1}\,\left(\bar{A}_{j}+(-1)^{m}\,\mathds{1}\right)\,\left\{\bar{A}_{j},\,\Gamma\right\}\,,~{}~{}$ (C.108)
@@
$\Gamma^{\prime}$ $=\frac{1}{2}\sum\limits_{m=0,1}\,(-1)^{m}\,P^{(m)}_{j}\,\sum\limits_{\mu=1}^{3}\,\hat{\alpha}_{j,\mu}\,\left\{\sigma^{\mu}_{j},\,\sigma^{\nu}_{j}\right\}$
$=\sum\limits_{m=0,1}\,(-1)^{m}\,P^{(m)}_{j}\,\hat{\alpha}_{j,\nu}=\operatorname*{tr}\limits_{j}\,\left[\,\bar{A}_{j}\,\sigma^{\nu}_{j}\,\right]\,\bar{A}_{j}\,,~{}~{}$ (C.109)
@@
In other words, if we evolve any logical operator $\Gamma$ (or Pauli-string observable of interest) under a dilated unitary channel $\mathcal{M}$ corresponding to projective measurement of a single-qubit observable $A_{j}$ without outcome-dependent recovery operations, then the Heisenberg-evolved operator $\Gamma^{\prime}$ (C.107) realizes one of the following:
@@
    If the operator $\Gamma$ and $A_{j}$ share no support, then $\Gamma^{\prime}=\Gamma$, meaning that $\mathcal{M}$ acts trivially.
@@
    If the involutory part of $A_{j}$ and the Pauli content of $\Gamma$ on qubit $j$ are both $\sigma^{\mu}_{j}$, then $\Gamma^{\prime}=\Gamma$, so that $\mathcal{M}$ acts trivially.
@@
    If the involutory part of $A_{j}$ and the Pauli content of $\Gamma$ on qubit $j$ correspond to distinct Pauli operators, then $\Gamma^{\prime}=0$, meaning that teleportation fails because the logical operator is destroyed.
@@
    If the involutory part of $A_{j}$ realizes a superposition of Pauli operator (4), then $\Gamma^{\prime}$ is no longer involutory nor unitary, and thus cannot reproduce the logical operator algebra (1) acting on the initial state, which is required by Prop. 1.
@@ Appendix D Proof of Proposition 6
$\rho_{\rm L}\equiv\operatorname*{tr}_{F^{c}}\left[\mathcal{W}\,\rho_{0}\otimes\lvert\boldsymbol{0}\rangle\langle\boldsymbol{0}\rvert_{\rm ss}\,\mathcal{W}^{\dagger}\right]\,,~{}~{}$ (D.110)
is a pure state, and separable with respect to the partition $F|F^{c}$. In particular, $\lvert F\rvert=k$ is the number of logical qubits, and $\rho_{\rm L}$ (D.110) is guaranteed to be unitarily equivalent to the pure state $\bigotimes_{n=1}^{k}\lvert\psi_{n}\rangle\langle\psi_{n}\rvert_{n}$. Because $F^{c}$ includes all Stinespring qubits, $\rho_{\rm L}$ (D.110) is implicitly averaged over the outcomes of all measurements in $\mathcal{W}$.
Additionally, the logical reduced density matrix $\rho_{\rm L}$ (D.110) after realizing the sequence of measurement outcomes $\boldsymbol{n}=(n_{1},\dots,n_{M})$ is given by Friedman et al. (2022b); Barberena and Friedman (2023); Friedman et al. (2022a)
$\rho^{(\boldsymbol{n})}_{\rm L}\equiv\frac{\operatorname*{tr}_{F^{c}}\left[\,\rho_{T}\,\widetilde{P}^{(\boldsymbol{n})}_{\rm ss}\,\right]}{\operatorname*{tr}_{\rm dil}\left[\,\rho_{T}\,\widetilde{P}^{(\boldsymbol{n})}_{\rm ss}\,\right]}\,,~{}~{}$ (D.111)
where the denominator is equal to the the probability $p_{\boldsymbol{n}}$ of the “outcome trajectory” $\boldsymbol{n}$.
Thus, the outcome-averaged logical reduced density matrix $\rho_{\rm L}$ (D.110) is automatically a convex sum over trajectories $\boldsymbol{n}$ of the reduced density matrix $\rho^{(\boldsymbol{n})}_{\rm L}$ (D.111) weighted by the scalar $p_{\boldsymbol{n}}$. However, such a density matrix is pure if and only if every term in the sum is proportional to $\rho_{\rm L}$ (D.110). Because $\rho_{\rm L}$ (D.110) and $\rho^{(\boldsymbol{n})}_{\rm L}$ (D.111) are proportional and normalized for all $\boldsymbol{n}$, we must have that $\rho^{(\boldsymbol{n})}_{\rm L}=\rho_{\rm L}$ for all $\boldsymbol{n}$. Hence, the output state $\rho_{\rm L}$ (D.110) is realized on all outcome trajectories. ∎
@@ Appendix E Proof of Lemma 7
Let $\mathcal{R}_{s}$ be a recovery channel (34) and $\Gamma$ to be be any logical operator. By construction (and Def. 5), both $\chi_{s}$ (34) and $\Gamma$ are Pauli-string operators (2). Evolving $\Gamma$ under $\mathcal{R}_{s}$ in the Heisenberg picture gives
@@
$\mathcal{R}^{\dagger}_{s}\,\Gamma\otimes\widetilde{\mathds{1}}\,\mathcal{R}_{s}$ $=\sum\limits_{m,n=0,1}\,\chi^{m}_{s}\,\Gamma\,\chi^{n}_{s}\otimes\widetilde{P}^{(m)}_{r_{s}}\,\widetilde{\mathds{1}}\,\widetilde{P}^{(n)}_{r_{s}}$
$=\sum\limits_{m=0,1}\,\chi^{m}_{s}\,\Gamma\,\chi^{m}_{s}\otimes\widetilde{P}^{(m)}_{r_{s}}$
$=\sum\limits_{m=0,1}\,\left(-1\right)^{m\,\lambda_{\Gamma,s}}\,\Gamma\otimes\widetilde{P}^{(m)}_{r_{s}}$
@@
$=\Gamma\otimes\left(\delta_{\lambda_{\Gamma,s},0}\,\widetilde{\mathds{1}}+\delta_{\lambda_{\Gamma,s},1}\,\widetilde{Z}_{r_{s}}\right)$
$=\Gamma\otimes\widetilde{Z}^{\lambda_{\Gamma,s}}_{r_{s}}\,,~{}~{}$ (E.112)
where $\widetilde{Z}_{r_{s}}$ is the product of $\widetilde{Z}_{j}$ for $j\in r_{s}$, where
$\chi_{s}\,\Gamma\,\chi_{s}=\left(-1\right)^{\lambda_{\Gamma,s}}\,\Gamma\,,~{}~{}$ (E.113)
defines $\lambda_{\Gamma,s}\in\{0,1\}$ for each $\Gamma$ and $\mathcal{R}_{s}$.
@@
Next, we prove Eq. 7 for a given Pauli string $\Gamma$. The result (E.112) establishes that evolving $\Gamma$ under any $\mathcal{R}$ multiplies $\Gamma$ by Stinespring operators $\widetilde{Z}_{j}$ for $j\in r_{s}$. Hence, any recovery channel $\mathcal{R}_{\Gamma}$ (7) attaches $\widetilde{Z}_{r}$ to $\Gamma$ provided that $\left\{\chi_{\Gamma},\,\Gamma\right\}=0$. The particular choice
$\chi_{\Gamma}=\prod\limits_{s=1}^{n}\chi^{\lambda_{\Gamma,s}}_{s}\,,~{}~{}$ (E.114)
@@
always fulfils this condition, where $\lambda_{\Gamma,s}\in\{0,1\}$ is defined in Eq. E.113 and $n$ is the number of Clifford recovery channels $\mathcal{R}_{s}$. However, any choice of Pauli string $\chi_{\Gamma}$ that anticommutes with $\Gamma$ leads to a channel $\mathcal{R}_{\Gamma}$ (7) that is equivalent to $\mathcal{R}$ in its action on $\Gamma$.
@@
Next, consider a set of logical operators $\boldsymbol{\Gamma}$, consisting of, e.g., all $X$- and $Z$-type logical operators for the $k$ logical qubits. Again, Eq. E.112 establishes that $\mathcal{R}$ acts by attaching (possibly different) $\widetilde{Z}_{j}$ operators to different logical operators $\Gamma\in\boldsymbol{\Gamma}$. However, by the previous result (7), $\mathcal{R}$ is equivalent in its action on each $\Gamma\in\boldsymbol{\Gamma}$ to some $\mathcal{R}_{\Gamma}$ (7) where $\chi_{\Gamma}$ is any Pauli string that anticommutes with $\Gamma$. Hence, we can write $\mathcal{R}$ as a product over $\Gamma\in\boldsymbol{\Gamma}$ of channels $\mathcal{R}_{\Gamma}$ (7) provided that $\chi_{\Gamma}$ anticommutes only with the logical operator $\Gamma$, and not its counterpart $\Gamma^{\prime}$, which acts on the same logical qubit. Since $\left\{\Gamma,\,\Gamma^{\prime}\right\}=\left[\Gamma^{\prime},\,\Gamma^{\prime}\right]=0$, the choice $\chi_{\Gamma}=\Gamma^{\prime}$ is sufficient to ensure that Eq. 37 is equivalent to $\mathcal{R}$ for all logical operators. We note that the decomposition (37) is not unique. ∎
@@ Appendix F Proof of Lemma 8
$\mathcal{R}_{j,k}$ $=\sum\limits_{n=0,1}\,\chi^{n}_{j,k}\otimes\widetilde{P}^{(n)}_{j}$ (F.115)
$\mathcal{M}_{j}\,$ $=\frac{1}{2}\sum\limits_{m=0,1}\left(\mathds{1}+(-1)^{m}\,\bar{A}^{\prime}_{j}\right)\otimes\widetilde{X}^{m}_{j}\,,~{}~{}$ (F.116)
where $\chi_{j,k}\in\mathbf{P}_{*}{\rm ph}$ is a physical Pauli string, $\bar{A}_{j}^{\prime}$ is the involutory part (4) of an operator of the form
$A_{j}^{\prime}=U_{j}\,A_{j}U^{\dagger}_{j}$ (F.117)
where $A_{j}$ is a single-qubit observable, and $U_{j}$ is a generic physical unitary. Note that Eq. F.117 captures modifications due to writing $\mathcal{W}^{\prime}$ in canonical form (31).
@@
$\Gamma(0)=\Gamma=\sigma^{\nu}_{f_{n}}\otimes\widetilde{\mathds{1}}_{\rm ss}\in\mathbf{P}_{\rm ph}\,,~{}~{}$ (F.118)
@@
which acts as a logical operator on the final state $\lvert\Psi_{T}\rangle$ (7). Evolving this operator in the Heisenberg picture, projecting onto the default Stinespring initial state $\lvert\boldsymbol{0}\rangle_{\rm ss}$, and averaging over outcomes leads to Eq. 1, i.e.,
@@
$\Gamma_{\rm ph}(T)=\operatorname*{tr}\limits_{\rm ss}\left[\,\mathcal{W}^{\dagger}\,\Gamma(0)\,\mathcal{W}\,\widetilde{P}^{(\boldsymbol{0})}_{\rm ss}\,\right]=\sigma^{\nu}_{i_{n}}\mathsf{S}_{n,\nu}\,,~{}~{}$ (F.119)
@@
which is the same logical operator $\Gamma$ acting on the initial state $\lvert\Psi_{0}\rangle$ (6), where $\mathsf{S}_{n,\nu}$ is an element of the initial-state stabilizer group $\operatorname*{stab}\left(\lvert\Psi_{0}\rangle\right)\subset\operatorname*{Aut}\left(*\right){\mathcal{H}_{\rm ph}}$ (3).
@@
Importantly, the naïve protocol $\mathcal{W}$ need not be in canonical form (31). Without loss of generality, we have omit any measurements without corresponding feedback and convert trivial recovery operations into physical unitaries, as described above. However, for convenience, we also pull all non-Clifford physical unitaries into $\mathcal{U}$, and if the combination of a measurement channel $\mathcal{M}_{j}$ is conditioned on a prior outcome in the naïve protocol $\mathcal{W}$, then we pull prior recovery channels through $\mathcal{M}_{j}$ as needed so that it realizes Eq. F.116. This is not necessary if $\mathcal{M}_{j}$ commutes with all prior recovery channels (in the Schrödinger picture) or if $A_{j}$ is a Pauli string.
@@
As a result, all operations after the first measurement channel (in the Schrödinger picture) correspond either to measurements or Clifford channels. Hence, in the Heisenberg picture, any logical operator $\Gamma(\tau_{j,k})$ prior to the recovery channel $\mathcal{R}_{s}$ is guaranteed to be a Pauli string, provided that no measurement channels have been encountered. Note that any number of recovery operations conditioned on various measurement outcomes can be factorized into the desired form (F.115), up to Clifford gates $U\in\mathbf{C}_{*}{\rm ph}$, by Def. 5 and Lemma 7.
@@
Suppose that $\mathcal{M}_{j}$ is the first measurement channel encountered in the Heisenberg picture. By Lemma 7,
$\mathcal{R}^{\dagger}_{j,k}\,\Gamma(\tau_{j,k})\,\mathcal{R}_{j,k}=\Gamma(\tau_{j,k})\,\widetilde{Z}^{\lambda_{\Gamma,j,k}}_{j}\,,~{}~{}$ (F.120)
where $\lambda_{\Gamma,j,k}$ is defined via
$\chi_{j,k}\,\Gamma(\tau_{j,k})\,\chi_{j,k}=\left(-1\right)^{\lambda_{\Gamma,j,k}}\,\Gamma(\tau_{j,k})\,,~{}~{}$ (E.113)
@@
so that the physical part of $\Gamma(\tau_{j,k})$ is unaffected by the recovery channel, while the Stinespring part is multiplied by the operator $\widetilde{Z}_{j}$ if and only if the recovery Pauli $\chi_{j,k}$ anticommutes with $\Gamma(\tau_{j,k})$ immediately prior to the recovery channel $\mathcal{R}_{j,k}$ (in the Heisenberg picture). If $\lambda_{\Gamma,j,k}=0$ in Eq. E.113 then Eq. F.120 is trivial, as claimed.

### v3

@@ Appendix C Proof of Proposition 5
At Heisenberg time $\tau=0$, the initial operator acts as $\Gamma\otimes\widetilde{\mathds{1}}^{\vphantom{*}}_{\rm ss}$ (i.e., trivially on all Stinespring registers). The outcome-averaged Heisenberg update to $\Gamma$ is given by
$\Gamma^{\prime}$ $=\operatorname*{tr}\limits_{{\rm ss},j}\,\left[\,\mathcal{M}^{\dagger}_{\left[A^{\,}_{j}\right]}\,\Gamma\otimes\widetilde{\mathds{1}}^{\vphantom{*}}_{j}\,\mathcal{M}^{\vphantom{\dagger}}_{\left[A^{\,}_{j}\right]}\,\widetilde{P}^{(0)}_{j}\,\right]\,,~{}~{}$ (C.107)
@@
$\Gamma^{\prime}$ $=\sum\limits_{m,n=0,1}\frac{1}{4}\left(\mathds{1}+(-1)^{m}\,\bar{A}^{\vphantom{\prime}}_{j}\right)\,\Gamma\,\left(\mathds{1}+(-1)^{m}\,\bar{A}^{\vphantom{\prime}}_{j}\right)$
×$~{}~{}~{}\quad~{}~{}\times~{}{\rm tr}\left[\,\widetilde{X}^{m}_{j}\,\widetilde{\mathds{1}}^{\vphantom{*}}_{j}\,\widetilde{X}^{n}_{j}\,\widetilde{P}^{(n)}_{j}\,\right]$
$=\frac{1}{4}\sum\limits_{m=0,1}\,\left(\Gamma+(-1)^{m}\,\left\{\bar{A}^{\,}_{j},\,\Gamma\right\}+\bar{A}^{\,}_{j}\,\Gamma\,\bar{A}^{\,}_{j}\right)$
$=\frac{1}{4}\sum\limits_{m=0,1}\,\left(\bar{A}^{\,}_{j}+(-1)^{m}\,\mathds{1}\right)\,\left\{\bar{A}^{\,}_{j},\,\Gamma\right\}\,,~{}~{}$ (C.108)
@@
$\Gamma^{\prime}$ $=\frac{1}{2}\sum\limits_{m=0,1}\,(-1)^{m}\,P^{(m)}_{j}\,\sum\limits_{\mu=1}^{3}\,\hat{\alpha}^{\vphantom{\prime}}_{j,\mu}\,\left\{\sigma^{\mu}_{j},\,\sigma^{\nu}_{j}\right\}$
$=\sum\limits_{m=0,1}\,(-1)^{m}\,P^{(m)}_{j}\,\hat{\alpha}^{\vphantom{\prime}}_{j,\nu}=\operatorname*{tr}\limits_{j}\,\left[\,\bar{A}^{\vphantom{\prime}}_{j}\,\sigma^{\nu}_{j}\,\right]\,\bar{A}^{\vphantom{\prime}}_{j}\,,~{}~{}$ (C.109)
@@
In other words, if we evolve any logical operator $\Gamma$ (or Pauli-string observable of interest) under a dilated unitary channel $\mathcal{M}$ corresponding to projective measurement of a single-qubit observable $A^{\,}_{j}$ without outcome-dependent recovery operations, then the Heisenberg-evolved operator $\Gamma^{\prime}$ (C.107) realizes one of the following:
@@
    If the operator $\Gamma$ and $A^{\,}_{j}$ share no support, then $\Gamma^{\prime}=\Gamma$, meaning that $\mathcal{M}$ acts trivially.
@@
    If the involutory part of $A^{\,}_{j}$ and the Pauli content of $\Gamma$ on qubit $j$ are both $\sigma^{\mu}_{j}$, then $\Gamma^{\prime}=\Gamma$, so that $\mathcal{M}$ acts trivially.
@@
    If the involutory part of $A^{\,}_{j}$ and the Pauli content of $\Gamma$ on qubit $j$ correspond to distinct Pauli operators, then $\Gamma^{\prime}=0$, meaning that teleportation fails because the logical operator is destroyed.
@@
    If the involutory part of $A^{\,}_{j}$ realizes a superposition of Pauli operator (33), then $\Gamma^{\prime}$ is no longer involutory nor unitary, and thus cannot reproduce the logical operator algebra (8) acting on the initial state, which is required by Prop. 1.
@@ Appendix D Proof of Proposition 6
$\rho^{\vphantom{\prime}}_{\rm L}\equiv\operatorname*{tr}_{F^{c}}\left[\mathcal{W}\,\rho^{\vphantom{\prime}}_{0}\otimes\lvert\boldsymbol{0}\rangle\langle\boldsymbol{0}\rvert^{\vphantom{\prime}}_{\rm ss}\,\mathcal{W}^{\dagger}\right]\,,~{}~{}$ (D.110)
is a pure state, and separable with respect to the partition $F|F^{c}$. In particular, $\lvert F\rvert=k$ is the number of logical qubits, and $\rho^{\,}_{\rm L}$ (D.110) is guaranteed to be unitarily equivalent to the pure state $\bigotimes_{n=1}^{k}\lvert\psi^{\,}_{n}\rangle\langle\psi^{\,}_{n}\rvert^{\,}_{n}$. Because $F^{c}$ includes all Stinespring qubits, $\rho^{\,}_{\rm L}$ (D.110) is implicitly averaged over the outcomes of all measurements in $\mathcal{W}$.
Additionally, the logical reduced density matrix $\rho^{\,}_{\rm L}$ (D.110) after realizing the sequence of measurement outcomes $\boldsymbol{n}=(n^{\,}_{1},\dots,n^{\,}_{M})$ is given by [47, 46, 42]
$\rho^{(\boldsymbol{n})}_{\rm L}\equiv\frac{\operatorname*{tr}_{F^{c}}\left[\,\rho^{\vphantom{\prime}}_{T}\,\widetilde{P}^{(\boldsymbol{n})}_{\rm ss}\,\right]}{\operatorname*{tr}_{\rm dil}\left[\,\rho^{\vphantom{\prime}}_{T}\,\widetilde{P}^{(\boldsymbol{n})}_{\rm ss}\,\right]}\,,~{}~{}$ (D.111)
where the denominator is equal to the the probability $p^{\,}_{\boldsymbol{n}}$ of the “outcome trajectory” $\boldsymbol{n}$.
Thus, the outcome-averaged logical reduced density matrix $\rho^{\,}_{\rm L}$ (D.110) is automatically a convex sum over trajectories $\boldsymbol{n}$ of the reduced density matrix $\rho^{(\boldsymbol{n})}_{\rm L}$ (D.111) weighted by the scalar $p^{\,}_{\boldsymbol{n}}$. However, such a density matrix is pure if and only if every term in the sum is proportional to $\rho^{\,}_{\rm L}$ (D.110). Because $\rho^{\,}_{\rm L}$ (D.110) and $\rho^{(\boldsymbol{n})}_{\rm L}$ (D.111) are proportional and normalized for all $\boldsymbol{n}$, we must have that $\rho^{(\boldsymbol{n})}_{\rm L}=\rho^{\vphantom{\prime}}_{\rm L}$ for all $\boldsymbol{n}$. Hence, the output state $\rho^{\,}_{\rm L}$ (D.110) is realized on all outcome trajectories. ∎
@@ Appendix E Proof of Lemma 7
Let $\mathcal{R}^{\,}_{s}$ be a recovery channel (34) and $\Gamma$ to be be any logical operator. By construction (and Def. 5), both $\chi^{\,}_{s}$ (34) and $\Gamma$ are Pauli-string operators (2). Evolving $\Gamma$ under $\mathcal{R}^{\,}_{s}$ in the Heisenberg picture gives
@@
$\mathcal{R}^{\dagger}_{s}\,\Gamma\otimes\widetilde{\mathds{1}}^{\vphantom{*}}\,\mathcal{R}^{\vphantom{\dagger}}_{s}$ $=\sum\limits_{m,n=0,1}\,\chi^{m}_{s}\,\Gamma\,\chi^{n}_{s}\otimes\widetilde{P}^{(m)}_{r^{\,}_{s}}\,\widetilde{\mathds{1}}^{\vphantom{*}}\,\widetilde{P}^{(n)}_{r^{\,}_{s}}$
$=\sum\limits_{m=0,1}\,\chi^{m}_{s}\,\Gamma\,\chi^{m}_{s}\otimes\widetilde{P}^{(m)}_{r^{\,}_{s}}$
$=\sum\limits_{m=0,1}\,\left(-1\right)^{m\,\lambda^{\,}_{\Gamma,s}}\,\Gamma\otimes\widetilde{P}^{(m)}_{r^{\,}_{s}}$
@@
$=\Gamma\otimes\left(\delta^{\,}_{\lambda^{\,}_{\Gamma,s},0}\,\widetilde{\mathds{1}}^{\vphantom{*}}+\delta^{\,}_{\lambda^{\,}_{\Gamma,s},1}\,\widetilde{Z}^{\vphantom{*}}_{r^{\,}_{s}}\right)$
$=\Gamma\otimes\widetilde{Z}^{\lambda^{\,}_{\Gamma,s}}_{r^{\,}_{s}}\,,~{}~{}$ (E.112)
where $\widetilde{Z}^{\vphantom{*}}_{r^{\,}_{s}}$ is the product of $\widetilde{Z}^{\vphantom{*}}_{j}$ for $j\in r^{\,}_{s}$, where
$\chi^{\vphantom{\dagger}}_{s}\,\Gamma\,\chi^{\vphantom{\dagger}}_{s}=\left(-1\right)^{\lambda^{\vphantom{\prime}}_{\Gamma,s}}\,\Gamma\,,~{}~{}$ (E.113)
defines $\lambda^{\vphantom{\prime}}_{\Gamma,s}\in\{0,1\}$ for each $\Gamma$ and $\mathcal{R}^{\,}_{s}$.
@@
Next, we prove Eq. 38 for a given Pauli string $\Gamma$. The result (E.112) establishes that evolving $\Gamma$ under any $\mathcal{R}$ multiplies $\Gamma$ by Stinespring operators $\widetilde{Z}^{\vphantom{*}}_{j}$ for $j\in r^{\,}_{s}$. Hence, any recovery channel $\mathcal{R}^{\,}_{\Gamma}$ (38) attaches $\widetilde{Z}^{\vphantom{*}}_{r}$ to $\Gamma$ provided that $\left\{\chi^{\,}_{\Gamma},\,\Gamma\right\}=0$. The particular choice
$\chi^{\vphantom{\dagger}}_{\Gamma}=\prod\limits_{s=1}^{n}\chi^{\lambda^{\vphantom{\prime}}_{\Gamma,s}}_{s}\,,~{}~{}$ (E.114)
@@
always fulfils this condition, where $\lambda^{\vphantom{\prime}}_{\Gamma,s}\in\{0,1\}$ is defined in Eq. E.113 and $n$ is the number of Clifford recovery channels $\mathcal{R}^{\,}_{s}$. However, any choice of Pauli string $\chi^{\,}_{\Gamma}$ that anticommutes with $\Gamma$ leads to a channel $\mathcal{R}^{\,}_{\Gamma}$ (38) that is equivalent to $\mathcal{R}$ in its action on $\Gamma$.
@@
Next, consider a set of logical operators $\boldsymbol{\Gamma}$, consisting of, e.g., all $X^{\vphantom{*}}$- and $Z^{\vphantom{*}}$-type logical operators for the $k$ logical qubits. Again, Eq. E.112 establishes that $\mathcal{R}$ acts by attaching (possibly different) $\widetilde{Z}^{\vphantom{*}}_{j}$ operators to different logical operators $\Gamma\in\boldsymbol{\Gamma}$. However, by the previous result (38), $\mathcal{R}$ is equivalent in its action on each $\Gamma\in\boldsymbol{\Gamma}$ to some $\mathcal{R}^{\,}_{\Gamma}$ (38) where $\chi^{\,}_{\Gamma}$ is any Pauli string that anticommutes with $\Gamma$. Hence, we can write $\mathcal{R}$ as a product over $\Gamma\in\boldsymbol{\Gamma}$ of channels $\mathcal{R}^{\,}_{\Gamma}$ (38) provided that $\chi^{\,}_{\Gamma}$ anticommutes only with the logical operator $\Gamma$, and not its counterpart $\Gamma^{\prime}$, which acts on the same logical qubit. Since $\left\{\Gamma,\,\Gamma^{\prime}\right\}=\left[\Gamma^{\prime},\,\Gamma^{\prime}\right]=0$, the choice $\chi^{\,}_{\Gamma}=\Gamma^{\prime}$ is sufficient to ensure that Eq. 37 is equivalent to $\mathcal{R}$ for all logical operators. We note that the decomposition (37) is not unique. ∎
@@ Appendix F Proof of Lemma 8
$\mathcal{R}^{\vphantom{\dagger}}_{j,k}$ $=\sum\limits_{n=0,1}\,\chi^{n}_{j,k}\otimes\widetilde{P}^{(n)}_{j}$ (F.115)
$\mathcal{M}^{\vphantom{\dagger}}_{j}\,$ $=\frac{1}{2}\sum\limits_{m=0,1}\left(\mathds{1}+(-1)^{m}\,\bar{A}^{\prime}_{j}\right)\otimes\widetilde{X}^{m}_{j}\,,~{}~{}$ (F.116)
where $\chi^{\,}_{j,k}\in\mathbf{P}^{\vphantom{*}}_{*}{\rm ph}$ is a physical Pauli string, $\bar{A}_{j}^{\prime}$ is the involutory part (33) of an operator of the form
$A_{j}^{\prime}=U^{\vphantom{\dagger}}_{j}\,A^{\,}_{j}U^{\dagger}_{j}$ (F.117)
where $A^{\,}_{j}$ is a single-qubit observable, and $U^{\,}_{j}$ is a generic physical unitary. Note that Eq. F.117 captures modifications due to writing $\mathcal{W}^{\prime}$ in canonical form (31).
@@
$\Gamma(0)=\Gamma=\sigma^{\nu}_{f_{n}}\otimes\widetilde{\mathds{1}}^{\vphantom{*}}_{\rm ss}\in\mathbf{P}^{\vphantom{*}}_{\rm ph}\,,~{}~{}$ (F.118)
@@
which acts as a logical operator on the final state $\lvert\Psi^{\,}_{T}\rangle$ (7). Evolving this operator in the Heisenberg picture, projecting onto the default Stinespring initial state $\lvert\boldsymbol{0}\rangle^{\,}_{\rm ss}$, and averaging over outcomes leads to Eq. 8, i.e.,
@@
$\Gamma^{\vphantom{\prime}}_{\rm ph}(T)=\operatorname*{tr}\limits_{\rm ss}\left[\,\mathcal{W}^{\dagger}\,\Gamma(0)\,\mathcal{W}\,\widetilde{P}^{(\boldsymbol{0})}_{\rm ss}\,\right]=\sigma^{\nu}_{i_{n}}\mathsf{S}^{\vphantom{\dagger}}_{n,\nu}\,,~{}~{}$ (F.119)
@@
which is the same logical operator $\Gamma$ acting on the initial state $\lvert\Psi^{\,}_{0}\rangle$ (6), where $\mathsf{S}^{\,}_{n,\nu}$ is an element of the initial-state stabilizer group $\operatorname*{stab}\left(\lvert\Psi^{\,}_{0}\rangle\right)\subset\operatorname*{Aut}\left(*\right){\mathcal{H}^{\,}_{\rm ph}}$ (3).
@@
Importantly, the naïve protocol $\mathcal{W}$ need not be in canonical form (31). Without loss of generality, we have omit any measurements without corresponding feedback and convert trivial recovery operations into physical unitaries, as described above. However, for convenience, we also pull all non-Clifford physical unitaries into $\mathcal{U}$, and if the combination of a measurement channel $\mathcal{M}^{\,}_{j}$ is conditioned on a prior outcome in the naïve protocol $\mathcal{W}$, then we pull prior recovery channels through $\mathcal{M}^{\,}_{j}$ as needed so that it realizes Eq. F.116. This is not necessary if $\mathcal{M}^{\,}_{j}$ commutes with all prior recovery channels (in the Schrödinger picture) or if $A^{\,}_{j}$ is a Pauli string.
@@
As a result, all operations after the first measurement channel (in the Schrödinger picture) correspond either to measurements or Clifford channels. Hence, in the Heisenberg picture, any logical operator $\Gamma(\tau^{\,}_{j,k})$ prior to the recovery channel $\mathcal{R}^{\,}_{s}$ is guaranteed to be a Pauli string, provided that no measurement channels have been encountered. Note that any number of recovery operations conditioned on various measurement outcomes can be factorized into the desired form (F.115), up to Clifford gates $U\in\mathbf{C}^{\vphantom{*}}_{*}{\rm ph}$, by Def. 5 and Lemma 7.
@@
Suppose that $\mathcal{M}^{\,}_{j}$ is the first measurement channel encountered in the Heisenberg picture. By Lemma 7,
$\mathcal{R}^{\dagger}_{j,k}\,\Gamma(\tau^{\vphantom{\prime}}_{j,k})\,\mathcal{R}^{\vphantom{\dagger}}_{j,k}=\Gamma(\tau^{\vphantom{\prime}}_{j,k})\,\widetilde{Z}^{\lambda^{\vphantom{\prime}}_{\Gamma,j,k}}_{j}\,,~{}~{}$ (F.120)
where $\lambda^{\,}_{\Gamma,j,k}$ is defined via
$\chi^{\vphantom{\dagger}}_{j,k}\,\Gamma(\tau^{\vphantom{\prime}}_{j,k})\,\chi^{\vphantom{\dagger}}_{j,k}=\left(-1\right)^{\lambda^{\vphantom{\prime}}_{\Gamma,j,k}}\,\Gamma(\tau^{\vphantom{\prime}}_{j,k})\,,~{}~{}$ (E.113)
@@
so that the physical part of $\Gamma(\tau^{\,}_{j,k})$ is unaffected by the recovery channel, while the Stinespring part is multiplied by the operator $\widetilde{Z}^{\vphantom{*}}_{j}$ if and only if the recovery Pauli $\chi^{\,}_{j,k}$ anticommutes with $\Gamma(\tau^{\,}_{j,k})$ immediately prior to the recovery channel $\mathcal{R}^{\,}_{j,k}$ (in the Heisenberg picture). If $\lambda^{\,}_{\Gamma,j,k}=0$ in Eq. E.113 then Eq. F.120 is trivial, as claimed.